OpenStudy (anonymous):

Suppose that n is an integer such that n/3 is 12 greater than n/2 . Which of the following is the value of n ?

OpenStudy (anonymous):

n/3=n/2+12 n/3-n/2=12 2n/6-3n/6=12 -n/6=12 -n=72 n=72

OpenStudy (anonymous):

thankz